# Sharp error bounds for edge-element discretisations of the high-frequency Maxwell equations

### v2

@@ Abstract
We prove sharp wavenumber-explicit error bounds for first- or second-type-Nédélec-element (a.k.a. edge-element) conforming discretisations, of arbitrary (fixed) order, of the variable-coefficient time-harmonic Maxwell equations posed in a bounded domain with perfect electric conductor (PEC) boundary conditions. The PDE coefficients are allowed to be piecewise regular and complex-valued; this set-up therefore includes scattering from a PEC obstacle and/or variable real-valued coefficients, with the radiation condition approximated by a perfectly matched layer (PML).
@@
The results of this paper are the first preasymptotic error bounds for the time-harmonic Maxwell equations using first-type Nédélec elements or higher-than-lowest-order second-type Nédélec elements. Furthermore, they are the first wavenumber-explicit results, even in the asymptotic regime, for Maxwell scattering problems with a non-empty scatterer.
@@ 1.1 Statement of the main result
with wavenumber $k$, posed in a bounded domain $\Omega\subset\mathbb{R}^{3}$ with outward-pointing unit normal vector $n$ and diameter $L$, where $E\in H_{0}({\rm curl}\,,\Omega)$ (i.e., $E\in H({\rm curl}\,,\Omega)$ with $E\times n=0$ on $\partial\Omega$), the data $f\in(H_{0}({\rm curl}\,,\Omega))^{*}$, and the coefficients $\mu$ and $\epsilon$ (the relative permeability and relative permittivity, respectively) satisfy $\operatorname{Re}\mu,\operatorname{Re}\epsilon\geq c>0$ (in the sense of quadratic forms) in $\Omega$. We are interested in this problem when $kL\gg 1$, i.e., the high-frequency regime.
@@
We study approximations to the solution of (1.1) using the $h$-version of the finite-element method ($h$-FEM), where accuracy is increased by decreasing the meshwidth $h$ while keeping the polynomial degree $p$ constant, and the (conforming) approximation space consists of the first type of Nédélec finite elements [57], whose definition is recapped in §10.2 below 111 Recall that Nédélec elements are often called edge elements because at the lowest order basis functions and degrees of freedom are associated with the edges of the mesh; at higher order the geometrical identification of basis functions and degrees of freedom is more complicated. ; note that we choose the convention that $p=1$ corresponds to lowest-order Nédélec elements. Since the second type of Nédélec finite elements [58], [56, §8.2], [23, §15.5.1] contains the first type, and our results depend only on best-approximation properties of the space, our results also hold for second-type Nédélec finite elements.
@@
This definition implies that if $\Omega$ is $C^{\ell}$ with respect to a partition, then $\partial\Omega$ is $C^{\ell}$ (since $\partial\Omega=\Gamma_{i,j}$ for some $i,j$).
@@
Suppose that Assumption 1.2 holds for an integer $m\geq 1$. Let $K\subset(0,\infty)$ be the set of $k$ such that, given $f\in(H_{0}({\rm curl}\,,\Omega))^{*}$, there exists a unique solution $E$ to (1.1); let $C_{\rm sol}$ then be the $L^{2}(\Omega)\to L^{2}(\Omega)$ norm of the solution operator $f\mapsto E$.
Let ${\cal H}_{h}\subset H_{0}({\rm curl}\,,\Omega)$ be the space of first- or second-type Nédélec finite-elements of degree $p\geq 0$ on a (curved) mesh satisfying Assumption 10.1 below, with maximal element size $h$.
If $p\leq m$ then, given $k_{0}>0$, there exist $C_{1},C_{2}>0$ such that, for all $k\in K$ with $k\geq k_{0}$, if
@@
then the Galerkin solution $E_{h}$ exists, is unique, and satisfies
@@
Furthermore, given $C_{\rm{osc}}>0$ there exists $C_{3}>0$ such that if the data $f$ is $k$-oscillatory with constant $C_{\rm{osc}}$ and regularity index $m$ (in the sense of Definition 2.7 below), then
@@
Theorem 1.3 involves the $L^{2}(\Omega)\to L^{2}(\Omega)$ norm of the solution operator $f\mapsto E$, denoted by $C_{\rm sol}$. We note the following: (i) With the definition (1.2) of the norm $\|\cdot\|_{H_{k}({\rm curl}\,,\Omega)}$, the $k$-dependence of $C_{\rm sol}$ is the same as the $L^{2}(\Omega)\to H_{k}({\rm curl}\,,\Omega)$ norm of the solution operator. (ii) When $\mu$ and $\epsilon$ are both constant multiples of the identity in part of the domain, $C_{\rm sol}\geq CkL$ – this can be proved by considering data that is a cut-off function multiplied by a plane wave; see, e.g., [12, §1.4.1], [53, Example 3.4]. (iii) When $\epsilon$ and $\mu$ are both real, the problem is self adjoint, and $C_{\rm sol}$ is given in terms of the distance from $k^{2}$ to the spectrum [15, §2.3, §3]. (iv) [30, Theorem 1.6] proved that the norm of the solution operator of the Helmholtz PML problem is bounded by the norm of the solution operator of the corresponding Helmholtz scattering problem; we expect that the same result holds for the Maxwell PML problem – indeed, this result was recently proved for the case of no scatterer and Cartesian PML in [21, Lemma 10]. If so, then $C_{\rm sol}\leq CkL$ when the problem is nontrapping; see [12, Theorem 1.1] (for certain nontrapping coefficients) and [71, §2] (for a nontrapping PEC obstacle).
@@ 1.2 The context and novelty of the main result
The error bounds in the asymptotic regime rely on the fact that, since the Helmholtz adjoint solution operator is compact as a map from $L^{2}$ to $H^{1}$, the $L^{2}$ norm of the error is smaller than the $H^{1}$ norm by the Aubin–Nitsche lemma (see, e.g., [18, Theorem 19.1]). Indeed, Galerkin orthogonality $\langle P(u-u_{h}),v_{h}\rangle=0$ for all finite-element functions $v_{h}$ implies that, with $\Pi_{h}$ the orthogonal projection onto the finite-element space,
@@
The natural bounds in the preasymptotic regime (i.e., the Helmholtz analogues of (1.4) and (1.5) above) were proved in [33] for Helmholtz problems satisfying only a Gårding inequality and an elliptic-regularity shift, following earlier work by [70, 72, 22, 5, 61, 11]. Central to this earlier work was the elliptic projection argument [26, 27], which used that the Helmholtz operator is coercive if a sufficiently large multiple of the identity is added. The key insight in [33] is that, in fact, this coercivity can be achieved by adding a smoothing operator, defined in terms of eigenfunctions of the real part of the Helmholtz operator (a Maxwell analogue of this is Lemma 7.11 below).
@@
If $S$ is smoothing operator such that $P+S$ is coercive (with coercivity constant independent of $k$) and $P$ satisfies the natural assumptions for elliptic regularity, then $(P^{*}+S)^{-1}$ has the regularity shift associated with $(P^{*})^{-1}$, but its norm is bounded independent of $k$. Furthermore, $(P^{*}+S)^{-1}S(P^{*})^{-1}$ is smoothing, with norm bounded by the norm of $(P^{*})^{-1}$.
@@
  • The papers [52, 53] show that the $hp$-FEM applied to (1.1) with constant $\mu$ and $\epsilon$ and analytic boundary does not suffer from the pollution effect if $p\geq C_{1}\log(kL)$ for any $C_{1}>0$, $hk/p\leq C_{2}$ for sufficiently small $C_{2}>0$, and $C_{\rm sol}\leq C_{3}(kL)^{N}$ for some $C_{3},N>0$. The results in [53] for impedance boundary conditions contain the result that, when $p$ is constant, the Galerkin solution is quasioptimal (with constant independent of $k$) when $(kh)^{p+1}C_{\rm sol}$ is sufficiently small (i.e., one power of $kh$ away from the optimal result); see [53, Proof of Lemma 9.5] (and note that $p=0$ corresponds to the lowest-order elements in [53], instead of $p=1$ here). The fixed-$p$ results in [52] for when the radiation condition is realised exactly on $\partial\Omega$ are more restrictive; see [52, Remark 4.19]. These arguments essentially use a result equivalent to Lemma 6.1 below, and then prove approximation results about the adjoint solution operator to bound the second quantity in (6.1) (following the ideas introduced in the Helmholtz context in [50, 51, 24, 6]).
@@
  • Very recently, the natural preasymptotic error bounds were proved in [48, Theorem 4.2] when $p=1$ for (1.1) with constant $\mu$ and $\epsilon$ and an impedance boundary condition on $\partial\Omega$, and when the $h$-FEM is implemented using Nédélec elements of the second type. Recall that the second-type elements have better approximation properties in the $L^{2}$ norm than the first type (see, e.g., [56, §8.2]), with this fact crucially used in [48, Equation A.2]. The analogous error bounds for continuous interior-penalty methods were proved in [48, Theorem 5.2]. These results do not use the duality arguments described above; instead the crucial ingredient is a bound on the norm of the Galerkin solution in terms of the data; see [48, Theorem 4.1] and the discussion in [48, Remark 4.2].
    The results of [48] built on earlier work studying the same set up and proving the analogous result for other $h$-version FEMs, including interior-penalty discontinuous Galerkin methods [28, Theorem 6.1], a different continuous interior penalty method using second-type Nédélec elements [47, Theorem 4.6], and hybridizable discontinuous Galerkin methods [25, Theorem 4.7], [46, Remark 5.1].
@@ 1.3 Outline
§2 states the main result (i.e., Theorem 1.3) in abstract form (see Theorem 2.9 below). Using the material in §37, §8 proves Theorem 2.9. §9 recalls the regularity results of Weber [68]. §10 recalls the definition and properties of Nédélec finite elements. §11 proves Theorem 1.3. §A shows that the Maxwell PML problem falls into the class of Maxwell problems described in §1.1. §B recaps scaling arguments used to prove interpolation results for Nédélec elements on curved meshes.
@@ 2.1 Abstract framework and assumptions
${\cal H}$ and ${\cal V}$ are Hilbert spaces with ${\cal H}\subset{\cal V}$, ${\cal H}$ dense in ${\cal V}$, and norms ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\cal H}$ and ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{{\cal V}}$. $P:{\cal H}\to{\cal H}^{*}$ is bounded and such that $P=\mathcal{D}-\mathcal{E}$, where $\mathcal{E}:{\cal V}\to{\cal V}^{*}$ is bounded and there exists $C_{{\mathcal{E}}}^{\prime}>0$ such that
@@
Finally ${\rm Ker}\,\mathcal{D}^{*}={\rm Ker}\,\mathcal{D}$, and there exist $C_{1},C_{2}>0$ such that
@@
to see that this is indeed the square of a norm on ${\cal H}$, note that the right-hand side can be written as $\langle\operatorname{Re}{\cal A}v,v\rangle_{{\cal H}^{*}\times{\cal H}}:=\tfrac{1}{2}\langle({\cal A}+{\cal A}^{*})v,v\rangle_{{\cal H}^{*}\times{\cal H}}$ for ${\cal A}$ equal to $\mathcal{D}$ plus $C_{2}$ multiplied by the appropriate Riesz map ${\cal V}\to{\cal V}^{*}$ in the inner product corresponding to ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{{\cal V}}$.
@@
Let $\{u_{n}\}\in{\rm Ker}\,\mathcal{D}$ with $u_{n}\to u$ in ${\cal V}$. We need to show that $u\in{\rm Ker}\,\mathcal{D}$. By (2.1), $\|u_{n}\|_{{\cal H}}=\|u_{n}\|_{{\cal V}}$. Since $u_{n}$ is bounded in ${\cal V}$, $u_{n}$ is bounded in ${\cal H}$. Since ${\cal H}$ is a Hilbert space, by passing to a subsequence, we see that there exists $w\in{\cal H}$ such that $u_{n}\rightharpoonup w$ as $n\to\infty$. We now show that $w\in{\rm Ker}\,\mathcal{D}$. Let ${\cal R}:{\cal H}^{*}\to{\cal H}$ be the Riesz map such that $\langle a,b\rangle_{{\cal H}\times{\cal H}^{*}}=(a,{\cal R}b)_{\cal H}$ for all $a\in{\cal H},b\in{\cal H}^{*}$. Since $u_{n}\in{\rm Ker}\,\mathcal{D}$ for all $n$ and $u_{n}\rightharpoonup w$ as $n\to\infty$, for all $v\in{\cal H}$,
@@
Let ${\cal Z}^{0}={\cal V}$, ${\cal Z}^{1}={\cal H}$, and ${\cal Z}^{j}\subset{\cal Z}^{j-1}$ for $j=1,\ldots,m+1$, with ${\cal V}$ dense in $({\cal Z}^{j})^{*}$ for $j\geq 1$. Given $P$ satisfying Assumption 2.1, there exists $C>0$ such that the following is true.
@@
$\left\|\Pi_{1}u\right\|_{{\cal Z}^{j}}\leq C\Big{(}\left\|\Pi_{1}u\right\|_{{\cal V}}+\sup_{v\in{\cal H},\|\iota v\|_{({\cal Z}^{j-2})^{*}}=1}\big{|}\big{\langle}\mathsf{D}\Pi_{1}u,\Pi_{1}v\big{\rangle}_{{\cal H}^{*}\times{\cal H}}\big{|}\Big{)}\quad\text{ for all }u\in{\cal H},$ (2.10)
@@
$\left\|\mathsf{E}\right\|_{{\cal Z}^{j}\to{\cal Z}^{j}}\leq C.$ (2.11)
@@
$\big{\|}\Pi_{0}^{{\cal V}}u\big{\|}_{{\cal Z}^{j}}\leq C\Big{(}\big{\|}\Pi_{0}^{{\cal V}}\mathsf{E}\Pi_{0}^{{\cal V}}u\big{\|}_{{\cal Z}^{j}}+\big{\|}\Pi_{0}^{{\cal V}}u\big{\|}_{{\cal V}}\Big{)}\quad\text{ for all }u\in{\cal V}.$ (2.12)
@@
(ii) $f\in Z^{m-1}$ with $\mathop{{\rm div}}f=0$ and (2.20) holds with $m+1$ replaced by $m-1$.

### v3

@@ Abstract
We prove sharp wavenumber-explicit error bounds for first- or second-family-Nédélec-element (a.k.a. edge-element) conforming discretisations, of arbitrary (fixed) order, of the variable-coefficient time-harmonic Maxwell equations posed in a bounded domain with perfect electric conductor (PEC) boundary conditions. The PDE coefficients are allowed to be piecewise regular and complex-valued; this set-up therefore includes scattering from a PEC obstacle and/or variable real-valued coefficients, with the radiation condition approximated by a perfectly matched layer (PML).
@@
The results of this paper are the first preasymptotic error bounds for the time-harmonic Maxwell equations using first-family Nédélec elements or higher-than-lowest-order second-family Nédélec elements. Furthermore, they are the first wavenumber-explicit results, even in the asymptotic regime, for Maxwell scattering problems with a non-empty scatterer.
@@ 1.1 Statement of the main result
with wavenumber $k$, posed in a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{3}$ with outward-pointing unit normal vector $n$ and diameter $L$, where $E\in H_{0}({\rm curl}\,,\Omega)$ (i.e., $E\in H({\rm curl}\,,\Omega)$ with $E\times n=0$ on $\partial\Omega$), the data $f\in(H_{0}({\rm curl}\,,\Omega))^{*}$, and the coefficients $\mu$ and $\epsilon$ (the relative permeability and relative permittivity, respectively) satisfy $\operatorname{Re}\mu,\operatorname{Re}\epsilon\geq c>0$ (in the sense of quadratic forms) in $\Omega$. We are interested in this problem when $kL\gg 1$, i.e., the high-frequency regime.
@@
We study approximations to the solution of (1.1) using the $h$-version of the finite-element method ($h$-FEM), where accuracy is increased by decreasing the meshwidth $h$ while keeping the polynomial degree $p$ constant, and the (conforming) approximation space consists of the first family (also called the “first type” or “first kind”) of Nédélec finite elements [58], whose definition is recapped in §10.2 below  [^1] Recall that Nédélec elements are often called edge elements because at the lowest order basis functions and degrees of freedom are associated with the edges of the mesh; at higher order the geometrical identification of basis functions and degrees of freedom is more complicated. ; note that we choose the convention that $p=1$ corresponds to lowest-order Nédélec elements. Since the second family of Nédélec finite elements [59], [57, §8.2], [23, §15.5.1] contains the first family, and our results depend only on best-approximation properties of the space, our results also hold for second-family Nédélec finite elements.
@@
If (1.1) has a solution for every $f\in L^{2}(\Omega)$, we define
[…]
otherwise $C_{\rm sol}:=\infty$. This definition implies that $C_{\rm sol}=\infty$ if (1.1) does not have a unique solution for every $f\in L^{2}(\Omega)$. Recall that, with the norm convention (1.2), the $L^{2}(\Omega)\to H_{k}({\rm curl}\,,\Omega)$ and $(H_{k}({\rm curl}\,,\Omega))^{*}\to H_{k}({\rm curl}\,,\Omega)$ norms are then both bounded by a $k$-independent multiple of $1+C_{\rm sol}$.
@@
Figure 1.1: An example of $\Omega$ (shaded) with $\overline{\Omega}=\cup_{j=1}^{n}\overline{\Omega_{j}}$ satisfying Definition 1.1.
This definition implies that if $\Omega$ is $C^{\ell}$ with respect to a partition, then $\partial\Omega$ is $C^{\ell}$ (since $\partial\Omega=\Gamma_{i,j}$ for some $i,j$). Figure 1.1 shows an example of $\overline{\Omega}=\cup_{j=1}^{n}\overline{\Omega_{j}}$ satisfying Definition 1.1.
@@
Suppose that Assumption 1.2 holds for an integer $m\geq 1$. Given $1\leq p\leq m$ and $k_{0},C_{\rm osc}>0$ there exist $C_{1},C_{2},C_{3}>0$ such that the following holds.
Let ${\cal H}_{h}\subset H_{0}({\rm curl}\,,\Omega)$ be the space of first- or second-family Nédélec finite-elements of degree $p$ on a (curved) mesh satisfying Assumption 10.1 below, with maximal element size $h$.
For all $k\geq k_{0}$ and $h>0$ satisfying
@@
the Galerkin solution $E_{h}$ exists, is unique, and satisfies
@@
Furthermore, if the data $f$ is $k$-oscillatory with constant $C_{\rm{osc}}$ and regularity index $m$ (in the sense of Definition 2.7 below), then
@@
Note that if (1.1) does not have a unique solution for every $f\in L^{2}(\Omega)$ then $C_{\rm sol}=\infty$ and (1.4) is never satisfied.
@@
Theorem 1.3 involves $C_{\rm sol}$ defined in (1.3); i.e., the $L^{2}(\Omega)\to L^{2}(\Omega)$ norm of the solution operator $f\mapsto E$. We note the following: (i) With the definition (1.2) of the norm $\|\cdot\|_{H_{k}({\rm curl}\,,\Omega)}$, the $k$-dependence of $C_{\rm sol}$ is the same as the $L^{2}(\Omega)\to H_{k}({\rm curl}\,,\Omega)$ norm of the solution operator. (ii) When $\mu$ and $\epsilon$ are both constant multiples of the identity in part of the domain, $C_{\rm sol}\geq CkL$ – this can be proved by considering data that is a cut-off function multiplied by a plane wave; see, e.g., [12, §1.4.1], [54, Example 3.4]. (iii) When $\epsilon$ and $\mu$ are both real, the problem is self adjoint, and $C_{\rm sol}$ is given in terms of the distance from $k^{2}$ to the spectrum [15, §2.3, §3]. (iv) [30, Theorem 1.6] proved that the norm of the solution operator of the Helmholtz PML problem is bounded by the norm of the solution operator of the corresponding Helmholtz scattering problem; we expect that the same result holds for the Maxwell PML problem.  [^2] For the case of no scatterer and Cartesian PML, [21, Lemma 10] proved that $C_{\rm sol}\leq C(kL)^{2}$; i.e., a $kL$ loss from the expected estimate. If so, then $C_{\rm sol}\leq CkL$ when the problem is nontrapping; see [12, Theorem 1.1] (for certain nontrapping coefficients) and [72, §2] (for a nontrapping PEC obstacle).
@@ 1.2 The context and novelty of the main result
The error bounds in the asymptotic regime rely on the fact that, since the Helmholtz adjoint solution operator is compact as a map from $L^{2}$ to $H^{1}$, the $L^{2}$ norm of the error is asymptotically smaller than the $H^{1}$ norm by the Aubin–Nitsche lemma (see, e.g., [18, Theorem 19.1]). Indeed, with $P$ denoting the Helmholtz operator, Galerkin orthogonality $\langle P(u-u_{h}),v_{h}\rangle=0$ for all finite-element functions $v_{h}$ implies that, with $\Pi_{h}$ the orthogonal projection onto the finite-element space,
@@
The natural bounds in the preasymptotic regime (i.e., the Helmholtz analogues of (1.5) and (1.6) above) were proved in [33] for Helmholtz problems satisfying only a Gårding inequality and an elliptic-regularity shift, following earlier work by [71, 73, 22, 5, 62, 11]. Central to this earlier work was the elliptic projection argument [26, 27], which used that the Helmholtz operator is coercive if a sufficiently large multiple of the identity is added. The key insight in [33] is that, in fact, this coercivity can be achieved by adding a smoothing operator $S$, defined in terms of eigenfunctions of the real part of the Helmholtz operator (a Maxwell analogue of this is Lemma 7.11 below).
@@
If $S$ is a smoothing operator such that $P+S$ is coercive (with coercivity constant independent of $k$) and $P$ satisfies the natural assumptions for elliptic regularity, then $(P^{*}+S)^{-1}$ has the regularity shift associated with $(P^{*})^{-1}$, but its norm is bounded independent of $k$. Furthermore, $(P^{*}+S)^{-1}S(P^{*})^{-1}$ is smoothing, with norm bounded by the norm of $(P^{*})^{-1}$.
@@
  • The papers [53, 54] show that the $hp$-FEM applied to (1.1) with constant $\mu$ and $\epsilon$ and analytic boundary does not suffer from the pollution effect if $p\geq C_{1}\log(kL)$ for any $C_{1}>0$, $hk/p\leq C_{2}$ for sufficiently small $C_{2}>0$, and $C_{\rm sol}\leq C_{3}(kL)^{N}$ for some $C_{3},N>0$. Although these analyses are geared towards $p$ growing with $k$, the results in [54] for impedance boundary conditions contain the result that, when $p$ is constant, the Galerkin solution is quasioptimal (with constant independent of $k$) when $(kh)^{p-1}C_{\rm sol}$ is sufficiently small (i.e., one power of $kh$ away from the optimal result); see [54, Proof of Lemma 9.5] (and note that $p=0$ corresponds to the lowest-order elements in [54], instead of $p=1$ here). The fixed-$p$ results in [53] for when the radiation condition is realised exactly on $\partial\Omega$ are more restrictive; see [53, Remark 4.19]. These arguments essentially use a result equivalent to Lemma 6.1 below, and then prove approximation results about the adjoint solution operator to bound the second quantity in (6.1) (following the ideas introduced in the Helmholtz context in [51, 52, 24, 6]).
@@
  • Very recently, the natural preasymptotic error bounds were proved in [48, Theorem 4.2] when $p=1$ for (1.1) with constant $\mu$ and $\epsilon$ and an impedance boundary condition on $\partial\Omega$, and when the $h$-FEM is implemented using Nédélec elements of the second family. Recall that the second-family elements have better approximation properties in the $L^{2}$ norm than the first family (see, e.g., [57, §8.2]), with this fact crucially used in [48, Equation A.2]. The analogous error bounds for continuous interior-penalty methods were proved in [48, Theorem 5.2]. These results do not use the duality arguments described above; instead the crucial ingredient is a bound on the norm of the Galerkin solution in terms of the data; see [48, Theorem 4.1] and the discussion in [48, Remark 4.2].
    The results of [48] built on earlier work studying the same set up and proving the analogous result for other $h$-version FEMs, including interior-penalty discontinuous Galerkin methods [28, Theorem 6.1], a different continuous interior penalty method using second-family Nédélec elements [47, Theorem 4.6], and hybridizable discontinuous Galerkin methods [25, Theorem 4.7], [46, Remark 5.1].
    Finally, we note that, since the preprint of the present paper appeared, [49] extended the results of [48] to $p>1$ using ideas from the present paper/[33]. Indeed, [49, Theorem 4.3] proved the natural preasymptotic error bounds for $p\in\mathbb{Z}^{+}$ for (1.1) with constant $\mu$ and $\epsilon$ and an impedance boundary condition on $\partial\Omega$, and when the $h$-FEM is implemented using Nédélec elements of the second family.
@@ 1.3 Outline
§2 states the main result (i.e., Theorem 1.3) in abstract form (see Theorem 2.9 below). The proof of Theorem 2.9 is given in §8; this proof uses intermediate results proved in §37. The proof of Theorem 1.3 is given in §11, using Theorem 2.9 and the material in §9 (a recap of the regularity results of Weber [69]) and §10 (a recap of the definition and properties of Nédélec finite elements). §A shows that the Maxwell PML problem falls into the class of Maxwell problems described in §1.1. §B recaps scaling arguments used to prove interpolation results for Nédélec elements on curved meshes.
@@ 2.1 Abstract framework and assumptions
Given a Hilbert space $\cal V$}, let $\cal V$$^{*}$ denote the anti-dual space, and let $\langle\cdot,\cdot\rangle_{{\cal V}^{*}\times{\cal V}}$ be the duality pairing that is linear with respect to the first argument and anti-linear with respect to the second argument.
@@
[…]
$\|P\|_{{\cal H}\to{\cal H}^{*}}+\|\mathcal{E}\|_{{\cal V}\to{\cal V}^{*}}\leq C_{2}$
and
@@
In addition, $P=\mathcal{D}$-$\mathcal{E}$ where ${\rm Ker}\,\mathcal{D}^{*}={\rm Ker}\,\mathcal{D}$ and
@@
[…]
@@
[…]
@@
Let ${\cal Z}^{0}={\cal V}$, ${\cal Z}^{1}={\cal H}$, ${\cal Z}^{j}\subset{\cal Z}^{j-1}$ for $j=1,\ldots,m+1$, with $\cal V$$^{*}$ dense in $({\cal Z}^{j})^{*}$ for $j\geq 1$. Let ${C_{\rm reg}}>0$, $C_{1},C_{2},C_{{\mathcal{E}}}^{\prime}>0$, and let $P$ satisfy Assumption 2.1 with these $C_{1},C_{2},C_{{\mathcal{E}}}^{\prime}>0$.
@@
$\left\|\Pi_{1}u\right\|_{{\cal Z}^{j}}\leq{C_{\rm reg}}\Big(\left\|\Pi_{1}u\right\|_{{\cal V}}+\sup_{v\in{\cal H},\|\iota v\|_{({\cal Z}^{j-2})^{*}}=1}\big|\big\langle\mathsf{D}\Pi_{1}u,\Pi_{1}v\big\rangle_{{\cal H}^{*}\times{\cal H}}\big|\Big)\quad\text{ for all }u\in{\cal H},$ (2.10)
@@
$\left\|\mathsf{E}\right\|_{{\cal Z}^{j}\to{\cal Z}^{j}}\leq{C_{\rm reg}}.$ (2.11)
@@
$\big\|\Pi_{0}^{{\cal V}}u\big\|_{{\cal Z}^{j}}\leq{C_{\rm reg}}\Big(\big\|\Pi_{0}^{{\cal V}}\mathsf{E}\Pi_{0}^{{\cal V}}u\big\|_{{\cal Z}^{j}}+\big\|\Pi_{0}^{{\cal V}}u\big\|_{{\cal V}}\Big)\quad\text{ for all }u\in{\cal V}.$ (2.12)
@@
(ii) $f\in Z^{m-1}$ with $\mathop{{\rm div}}f=0$ and (2.20) holds with $m+1$ replaced by m1m-1.